# thanks: corresponding authorthanks: corresponding author

### v3

@@ 2.2.2 Heuristic confidence metrics from cluster statistics
We construct heuristic confidence metrics by leveraging cluster statistics obtained from clustering-based decoders, such as cluster sizes and aggregated error probabilities within clusters. Intuitively, smaller error clusters indicate better-localized error configurations in terms of the connectivity structure on the Tanner graph, resulting in less ambiguity during error inference. As the size and number of clusters increase, the portion of a logical operator’s support that lies outside the clusters will shrink, and therefore the probability of decoding failure (i.e., the probability that a logical error occurs due to an error chain existing outside the clusters) will increase; see Fig. 1(b).
@@
For a particular set of error mechanisms $\mathcal{E}$, clusters $\quantity{C_{i}}_{i}$ (such that $\forall i,C_{i}\subseteq\mathcal{E}_{i}$), and a positive real number $\alpha$, the cluster size $\alpha$-norm fraction is defined as
@@
For a particular set of error mechanisms $\mathcal{E}$, clusters $\quantity{C_{i}}_{i}$ (such that $\forall i,C_{i}\subseteq\mathcal{E}_{i}$), and a positive real number $\alpha$, the cluster LLR $\alpha$-norm fraction is defined as
@@
The set of error mechanisms $\mathcal{E}$ in the definitions may be the full error mechanism set of a DEM, or a subset when disjoint groups of mechanisms are decoded independently. For instance, in Calderbank–Shor–Steane (CSS) codes, if circuits involve only logical $Z$ resets and measurements, it suffices to take $\mathcal{E}$ as the set of $X$ errors to capture confidence information, provided that correlations between $X$ and $Z$ errors are ignored. This approach prevents errors unrelated to an observable from affecting its confidence evaluation.
@@ 2.2.3 Global post-selection strategy
Execute a clustering-based decoder and collect cluster size or LLR data (optionally restricted to error mechanisms involving the observables to be corrected). Given a method $m\in\quantity{\mathrm{size},\mathrm{LLR}}$, norm order $\alpha>0$, and cutoff threshold $c\in[0,1]$, accept the decoding result only when $Q_{m}^{(\alpha)}\leq c$.
@@
    No comparative decoding is required; a single decoder execution suffices to compute the metrics, while computing the logical gap requires $2^{k}$ executions for $k$ logical observables.
@@ 2.3 Real-time post-selection strategy
We extend our post-selection strategy to a real-time setting by integrating it with the sliding-window decoding framework for general QLDPC codes [42, 43, 44, 45, 32, 46]. Rather than collecting clusters statistics from global decoding outcomes, we rely only on data from a constant number of recent windows as shown in Fig. 1(c), enabling us to compute a cluster-confidence metric in real time and to triger early aborts when necessary.
@@
To integrate our post-selection strategy with the sliding-window framework, we employ a clustering-based decoder as the inner decoder. As described in Fig. 1(c), we track fractions of clusters overlapping with committed regions and evaluate a cluster norm fraction metric based on the $L$ most recent committed regions. Importantly, whether to abort is judged after decoding each window, potentially reducing the retry cost. The detailed strategy is as follows:
@@ 2.4.1 Global strategy analysis
Our primary focus is quantifying the trade-off between $p_{\mathrm{log}}$ and the abort rate $p_{\mathrm{abort}}$: How much can the logical error rate be reduced by aborting a certain fraction of low-confidence samples? For comparison, we consider two other easy-to-compute baseline metrics:
  • Correction weight: The log-likelihood weight of the decoder’s correction, as defined in Eq. (1).
  • Detector density: The fraction of violated detectors relative to the total detector count, which generalizes the “non-equilibrium magnetization” heuristic from Ref. [28].
Both baseline metrics have inverse correlation with decoder confidence. Notably, detector density offers exceptional computational efficiency as it can be computed before decoder execution, unlike other post-decoding confidence metrics. For surface codes, we additionally evaluate the logical gap from the MWPM decoder.
@@
Figure 2: Post-selection analysis of global decoding for rotated surface codes. (a) Logical error rates $p_{\mathrm{log}}$ are plotted against $p_{\mathrm{abort}}$ for Strategy 1 based on the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$, across different physical error rates $p\in\{0.001,0.003,0.005,0.01\}$ and code distances $d\in\{5,9,13\}$. The values of $p_{\mathrm{log}}$ without post-selection ($p_{\mathrm{abort}}=0$) are emphasized as filled circles. For $p=0.001$, an inset displaying the same data with a logarithmic scale on $p_{\mathrm{abort}}$ is included. Shaded regions indicate 95% confidence intervals. (b) Our strategy based on $Q_{\mathrm{LLR}}^{(2)}$ is compared with three other baseline strategies (the logical gap calculated by MWPM, correction weight, and detector density) for $d=13$ and $p=0.005$. (c) Various strategies are compared at a fixed abort rate $p_{\mathrm{abort}}=0.3$. We consider four norm orders $\alpha\in\{0.5,1,2,\infty\}$ (specified above or below the markers) for the cluster size and LLR norm fractions. (d) Required abort rates $p_{\mathrm{abort}}$ to achieve target values of $p_{\mathrm{log}}\in\{10^{-3},10^{-6}\}$ are plotted against $p$ when using the metric $Q_{\mathrm{LLR}}^{(2)}$.
@@
In Fig. 2(b), we benchmark our cluster-based approach against established methods for $(d,p)=(13,0.005)$, comparing $Q_{\mathrm{LLR}}^{(2)}$ with the logical gap (computed via MWPM), correction weight, and detector density. As expected, our method’s performance falls between the optimal logical gap strategy and the simpler baseline metrics.
@@
Figure 3: Post-selection analysis of global decoding for bivariate bicycle codes. Two variants of bivariate bicycle codes are considered: $[[144,12,12]]$ and $[[72,12,6]]$. Shaded regions represent 95% confidence intervals. (a) Any-observable logical error rates $p_{\mathrm{log}}$ are plotted against the abort rate $p_{\mathrm{abort}}$ at $p\in\{0.001,0.003,0.005\}$ for Strategy 1 based on $Q_{\mathrm{LLR}}^{(2)}$. For $p=0.001$, a log-scale inset is included. (b), (c) Various strategies are compared for the $[[144,12,12]]$ BB code at $p=0.003$. $p_{\mathrm{abort}}$ is fixed to 0.3 in (c). (d) Required $p_{\mathrm{abort}}$ to achieve target values of $p_{\mathrm{log}}\in\{10^{-3},10^{-6}\}$ are plotted against pp.
@@
Figure 3 presents the simulation results for two instances of BB codes: $[[72,12,6]]$ and $[[144,12,12]]$ [36]. (Here, $[[n,k,d]]$ characterizes a code with $n$ physical qubits, $k$ logial qubits, and the code distance $d$.) In Fig. 3(a), $p_{\mathrm{log}}$ is plotted against $p_{\mathrm{abort}}$ at $p\in\quantity{0.001,0.003,0.005}$, where the subplot for $p=0.001$ includes a log-scale inset. Various strategies are compared in Fig. 3(b) and (c) for the $[[144,12,12]]$ code at $p=0.003$, confirming that the performance of the cluster-based strategy surpasses the correction weight strategy by up to around two orders of magnitude and does not severely depend on the norm order $\alpha$ as long as $\alpha\geq 1$. The required abort rates to reach target values $p_{\mathrm{log}}\in\quantity{10^{-3},10^{-6}}$ are plotted against $p$ in Fig. 3(d).
Figure 4: Post-selection analysis of global decoding for a hypergraph product code. We consider a $[[225,9,6]]$ $(3,4)$-regular hypergraph product code, defined from the product of a 12-variable classical LDPC codes with itself. (a) Any-observable logical error rates $p_{\mathrm{log}}$ are plotted against the abort rate $p_{\mathrm{abort}}$ at $p=0.001$, comparing Strategy 1 based on $Q_{\mathrm{LLR}}^{(2)}$ with the baseline metrics. Shaded regions represent 95% confidence intervals. (b) Various strategies are compared at a fixed $p_{\mathrm{abort}}=0.3$.
@@
We evaluate our approach on a $(3,4)$-regular HGP code with parameters $[[225,9,6]]$, constructed by taking the product of a 12-variable classical LDPC code with itself, whose check matrix is given in the Methods section. This particular code offers practical advantages with its depth-8 syndrome extraction circuit [46], which we employ in our simulations. Figure 4 shows the results at $p=0.001$, where our cluster-based approach achieves approximately two orders of magnitude improvement in $p_{\mathrm{log}}$ compared to baseline methods, which is consistent with the performance gains observed for BB codes.
Additional simulation results are provided in the Supplementary Material, including comprehensive data for all parameter combinations (Supplementary Figures 1–3), a detailed examination of the influence of the norm order $\alpha$ (Supplementary Figure 4), and an evaluation of the alternative “conv-max-conf” setting, which treats BP-converged samples as having maximum decoding confidence (Supplementary Figure 5). The conv-max-conf setting avoids the need to modify BP+LSD to enforce LSD execution regardless of BP convergence, but is found to degrade post-selection performance in most cases.
@@ 2.4.2 Real-time strategy analysis
We now analyze the real-time post-selection strategy. To fairly assess the impact of “aborting mid-way”, we evaluate the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$ instead of $p_{\mathrm{abort}}$, by summing the total number of syndrome extraction rounds elapsed across all samples (including both aborted and accepted runs) and dividing by the number of accepted samples. In other words, $\overline{T}_{\mathrm{accepted}}$ represents the expected time cost required to obtain a single accepted run when immediately retrying after each aborted run. For comparison, global strategies yield $\overline{T}_{\mathrm{accepted}}=T/(1-p_{\mathrm{abort}})$.
@@
Our analysis focuses on $T$-round memory experiments using the surface code with distance $d=13$ and the $[[144,12,12]]$ bivariate bicycle (BB) code. We assume the same circuit-level noise model as the global strategy analysis with $p\in\{0.003,0.005\}$. The decoding is performed using the sliding window framework with window parameters $(W,F)=(5,1)$ for the surface code and $(W,F)=(3,1)$ for the BB code. In both cases, the BP+LSD decoder serves as the inner decoder. For post-selection, we employ the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$ as our confidence metric with the lookback window size $L\in\{1,2,3,5,7\}$. By varying the cutoff value for this metric, we obtain different trade-offs between $p_{\mathrm{log}}$ and the average time cost per accepted shot $\overline{T}_{\mathrm{accepted}}$.
@@
Figure 5 presents our main results for both codes with $T=d$, clearly exhibiting the trade-off relation between $p_{\mathrm{log}}$ and $\overline{T}_{\mathrm{accepted}}$. For comparison, we include two baseline results: the global post-selection strategy (dashed lines) and standard sliding-window decoding without post-selection ($\times$” marks). Notably, our real-time strategy achieves performance comparable to the global strategy in most cases, and specifically for the BB code at $p=0.005$, it even surpasses global post-selection performance by at most around 1.5 orders of magnitude in $p_{\mathrm{log}}$.
@@
Figure 5: Analysis of real-time post-selection via Strategy 2 for (a) the surface code with $d=13$ and (b) the $[[144,12,12]]$ bivariate bicycle code. Logical error rates $p_{\mathrm{log}}$ are plotted against the average time cost per accepted shots $\overline{T}_{\mathrm{accepted}}$, which represents the average time cost required to succeed when immediately retrying after each aborted attempt. The results are for the memory experiments with $T=d$ at $p\in\{0.003,0.005\}$, decoded with the $(5,1)$ or $(3,1)$ sliding window method. The LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$ is used for the metric and the parameter $L$ of the strategy varies across $\{1,2,3,5,7\}$. For comparison, the values for the global strategy (Strategy 1) and those for sliding window decoding without post-selection are presented additionally as dashed lines and $\times$’ marks, respectively. Shaded regions represent 95% confidence intervals.
@@
The real-time strategy can be particularly useful for circuits involving many syndrome extraction rounds, where the global strategy may become computationally prohibitive. To assess the practical viability of our approach in such scenarios, we investigate how its performance scales with the total number of rounds $T$. Figure 6(a) examines the scaling behavior by plotting both $p_{\mathrm{log}}/T$ and the acceptance rate $1-p_{\mathrm{abort}}$ (in a logarithmic scale) against $T$ for the $[[144,12,12]]$ BB code under various fixed cutoff values, including the scenario without post-selection as a baseline. (Note that we here plot the acceptance rate rather than $\overline{T}_{\mathrm{accepted}}$, as it directly reveals its decay over rounds and allows us to estimate the per-round abort rate.)
Two key observations emerge from Fig. 6(a): First, $p_{\mathrm{log}}/T$ remains nearly constant across different $T$ values even under post-selection, which indicates that logical errors accumulate linearly with the number of rounds, preserving the expected scaling behavior shown in the case without post-selection. Second, $\log(1-p_{\mathrm{abort}})$ exhibits clear linear dependence on $T$. This behavior can be interpreted as every single round has an equal per-round abort rate $p_{\mathrm{abort}}^{\mathrm{round}}$, which is related to $p_{\mathrm{abort}}$ as $p_{\mathrm{abort}}=1-(1-p_{\mathrm{abort}}^{\mathrm{round}})^{T}$ and thus can be estimated from the slope of the fit in the right panel. In Fig. 6(b), we plot the estimated values of $p_{\mathrm{log}}/T$ (at $T=24$) and $p_{\mathrm{abort}}^{\mathrm{round}}$ against the cutoff, clearly illustrating the trade-off relation between them.

### v4

@@ 2.2.2 Heuristic confidence metrics from cluster statistics
Figure 2: Overview of our heuristic cluster-based confidence metrics and post-selection strategies based on them. (a) After constructing valid clusters from a clustering-based decoder, cluster statistics are used to quantify decoding confidence. As the size and number of clusters grow, a larger portion of the logical operator’s support tends to overlap with clusters, thereby increasing the probability of logical failure. Based on this intuition, we define two metrics (the cluster size and LLR norm fractions) in Definitions 1 and 2, and introduce post-selection strategies that abort a trial when the metric value exceeds a chosen cutoff. (b) This approach is further extended to real-time decoding with the sliding-window framework. After decoding each window of $W$ rounds and committing its first $F$ rounds, clusters are constructed from committed errors within the latest LFLF rounds (where $L$ denotes the lookback window size), and a metric is evaluated based on them. If the value exceeds the cutoff, the trial is immediately aborted, thereby avoiding unnecessary computation.
We construct heuristic confidence metrics by leveraging cluster statistics obtained from clustering-based decoders, such as cluster sizes and aggregated error probabilities within clusters. Intuitively, smaller error clusters indicate better-localized error configurations in terms of the connectivity structure on the Tanner graph, resulting in less ambiguity during error inference. As the size and number of clusters increase, the portion of a logical operator’s support that lies outside the clusters will shrink, and therefore the probability of decoding failure (i.e., the probability that a logical error occurs due to an error chain existing outside the clusters) will increase; see Fig. 2(a).
@@
For a set of error mechanisms $\mathcal{E}$, disjoint clusters $\quantity{C_{i}}_{i}$ (such that $\forall i,C_{i}\subseteq\mathcal{E}$), and a positive real number $\alpha$, the cluster size $\alpha$-norm fraction is defined as
@@
For a set of error mechanisms $\mathcal{E}$, disjoint clusters $\quantity{C_{i}}_{i}$ (such that $\forall i,C_{i}\subseteq\mathcal{E}$), and a positive real number $\alpha$, the cluster LLR $\alpha$-norm fraction is defined as
@@
The set of error mechanisms $\mathcal{E}$ in the definitions may be the full error mechanism set of a DEM, or a subset when disjoint groups of mechanisms are decoded independently. For instance, in Calderbank–Shor–Steane (CSS) codes, if circuits involve only logical $Z$ resets and measurements, it suffices to take $\mathcal{E}$ as the set of $X$ errors to capture confidence information, provided that correlations between $X$ and $Z$ errors are ignored. This approach prevents errors unrelated to the logical Pauli correction under decoding from affecting its confidence evaluation.
@@ 2.2.3 Global post-selection strategy
Execute a clustering-based decoder and collect cluster size or LLR data (optionally restricted to a subset of error mechanisms that nontrivially affects the logical Pauli correction under decoding). Given a method $m\in\quantity{\mathrm{size},\mathrm{LLR}}$, norm order $\alpha>0$, and cutoff threshold $c\in[0,1]$, accept the decoding result only when $Q_{m}^{(\alpha)}\leq c$.
@@
    No comparative decoding is required; a single decoder execution suffices to compute the metrics, while computing the logical gap requires at most $4^{k}$ executions for $k$ logical qubits.
@@ 2.3 Real-time post-selection strategy
We extend our post-selection strategy to a real-time setting by integrating it with the sliding-window decoding framework for general QLDPC codes [40, 35, 4, 60, 39, 42]. Rather than collecting clusters statistics from global decoding outcomes, we rely only on data from a constant number of recent windows as shown in Fig. 2(b), enabling us to compute a cluster-confidence metric in real time and to triger early aborts when necessary.
@@
To integrate our post-selection strategy with the sliding-window framework, we employ a clustering-based decoder as the inner decoder. As described in Fig. 2(b), we track fractions of clusters overlapping with committed regions and evaluate a cluster norm fraction metric based on the $L$ most recent committed regions. Importantly, whether to abort is judged after decoding each window, potentially reducing the retry cost. The detailed strategy is as follows:
@@ 2.4.1 Global strategy analysis
Our primary focus is quantifying the trade-off between $p_{\mathrm{log}}$ and the abort rate $p_{\mathrm{abort}}$: How much can the logical error rate be reduced by aborting a certain fraction of low-confidence samples? Since no established confidence metric exists for general QLDPC codes beyond the logical gap via comparative decoding, we compare against two baseline metrics that represent fundamental soft outputs applicable to any decoder with minimal computational overhead:
  • Correction weight: The log-likelihood weight of the decoder’s correction, as defined in Eq. (1). This quantity is the cost function that (non-degenerate) maximum likelihood decoders including MWPM directly minimize, and is related to error likelihood via $\Pr(E)\propto\exp(-w(E))$. High correction weight indicates an atypical, noisy shot, suggesting lower confidence in the decoded result.
  • Detector density: The fraction of violated detectors relative to the total detector count. Like correction weight, detector density quantifies how noisy and atypical a shot is. Unlike other metrics, however, it can be computed before decoder execution, offering exceptional computational efficiency (though it is not strictly a decoder soft output). This metric is related to the annular-syndrome post-selection rule in Ref. [8] and extends the “non-equilibrium magnetization” heuristic from Ref. [29] (originally developed for code-capacity noise) to circuit-level noise, motivated by the insight that syndrome density serves as an order parameter in the associated statistical-mechanical spin model [29, 21].
Both heuristic baselines are expected to have inverse correlation with decoder confidence, which we confirm numerically.
Additionally, we compare with the logical gap, which requires comparative decoding across logical classes. For surface codes (with a single logical qubit), this is computationally tractable, thus we evaluate the logical gap using both MWPM and BP+LSD to assess whether the method extends beyond MWPM decoders. For BB and HGP codes where exhaustive comparative decoding requires enormous computational resources, we try approximating the logical gap (referred to as the “logical gap proxy”) by decoding within a subset of $N_{\mathrm{class}}$ classes composed of the standard BP+LSD correction and $N_{\mathrm{class}}-1$ other logical classes selected randomly. Specifically, we use $N_{\mathrm{class}}=24$ for the $[[144,12,12]]$ BB code and $N_{\mathrm{class}}=18$ for the $[[225,9,6]]$ HGP code. (We also test more sophisticated heuristics for the gap proxy based on the empirical logical error distribution, but none outperform uniform random sampling; see the Supplementary Material for details.)
@@
Figure 3: Post-selection analysis of global decoding for rotated surface codes. (a) Logical error rates $p_{\mathrm{log}}$ are plotted against $p_{\mathrm{abort}}$ for Strategy 1 based on the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$, across different physical error rates $p\in\{0.001,0.003,0.005,0.01\}$ and code distances $d\in\{5,9,13\}$. The values of $p_{\mathrm{log}}$ without post-selection ($p_{\mathrm{abort}}=0$) are emphasized as filled circles. For $p=0.001$, an inset displaying the same data with a logarithmic scale on $p_{\mathrm{abort}}$ is included. Shaded regions indicate 95% confidence intervals. (b) Our strategy based on $Q_{\mathrm{LLR}}^{(2)}$ is compared with four other strategies for $d=13$ and $p=0.005$: the logical gap calculated by MWPM, the logical gap calculated by BP+LSD, correction weight, and detector density. (c) Various strategies are compared at a fixed abort rate $p_{\mathrm{abort}}=0.3$. We consider four norm orders $\alpha\in\{0.5,1,2,\infty\}$ (specified above or below the markers) for the cluster size and LLR norm fractions. (d) Required abort rates $p_{\mathrm{abort}}$ to achieve target values of $p_{\mathrm{log}}\in\{10^{-3},10^{-6}\}$ are plotted against $p$ when using the metric $Q_{\mathrm{LLR}}^{(2)}$.
@@
In Fig. 3(b), we benchmark our cluster-based approach against established methods for $(d,p)=(13,0.005)$, comparing $Q_{\mathrm{LLR}}^{(2)}$ with the logical gap computed via both MWPM and BP+LSD, correction weight, and detector density. As expected, our method’s performance falls between the near-optimal MWPM logical gap strategy and the simpler baseline metrics. Notably, the BP+LSD logical gap exhibits significantly degraded performance compared to MWPM (achieving only $p_{\mathrm{log}}\sim 10^{-5}$ versus $\sim 10^{-8}$ at $p_{\mathrm{abort}}=0.3$) despite using the same comparative decoding framework. We attribute this degradation to BP+LSD not guaranteeing the minimum-weight correction within each logical class, which can cause the correction weights to misrepresent the true likelihood of each class (as discussed in the “Logical gap for quantifying decoding confidence” section).
@@
Figure 4: Post-selection analysis of global decoding for bivariate bicycle codes. Two variants of bivariate bicycle codes are considered: $[[144,12,12]]$ and $[[72,12,6]]$. Shaded regions represent 95% confidence intervals. (a) Any-observable logical error rates $p_{\mathrm{log}}$ are plotted against the abort rate $p_{\mathrm{abort}}$ at $p\in\{0.001,0.003,0.005\}$ for Strategy 1 based on $Q_{\mathrm{LLR}}^{(2)}$. For $p=0.001$, a log-scale inset is included. (b), (c) Various strategies are compared for the $[[144,12,12]]$ BB code at $p=0.003$, including a logical gap proxy (“random” heuristic) calculated from a subset of logical classes composed of the standard BP+LSD correction and $N_{\text{class}}-1=23$ other classes selected randomly. $p_{\mathrm{abort}}$ is fixed to 0.3 in (c). (d) Required $p_{\mathrm{abort}}$ to achieve target values of $p_{\mathrm{log}}\in\{10^{-3},10^{-6}\}$ are plotted against pp.
@@
Figure 4 presents the simulation results for two instances of BB codes: $[[72,12,6]]$ and $[[144,12,12]]$ [11]. (Here, $[[n,k,d]]$ characterizes a code with $n$ physical qubits, $k$ logical qubits, and the code distance $d$.) In Fig. 4(a), $p_{\mathrm{log}}$ is plotted against $p_{\mathrm{abort}}$ at $p\in\quantity{0.001,0.003,0.005}$, where the subplot for $p=0.001$ includes a log-scale inset. Various strategies are compared in Fig. 4(b) and (c) for the $[[144,12,12]]$ code at $p=0.003$, confirming that the performance of the cluster-based strategy surpasses the correction weight strategy by up to around two orders of magnitude and does not severely depend on the norm order $\alpha$ as long as $\alpha\geq 1$.
Additionally, the gap proxy method with $N_{\mathrm{class}}=24$ fails to outperform even the correction weight despite the (at least) 24-fold increase in computational cost compared to standard decoding. This may be attributed to two factors: (i) The logical gap method itself is less effective for BP+LSD than for MWPM, as demonstrated by the surface code results above, and (ii) approximating the gap using only a subset of randomly sampled logical classes introduces additional estimation error, further degrading performance.
The required abort rates to reach target values $p_{\mathrm{log}}\in\quantity{10^{-3},10^{-6}}$ are plotted against $p$ in Fig. 4(d).
Figure 5: Post-selection analysis of global decoding for a hypergraph product code. We consider a $[[225,9,6]]$ $(3,4)$-regular hypergraph product code, defined from the product of a 12-variable classical LDPC codes with itself. (a) Any-observable logical error rates $p_{\mathrm{log}}$ are plotted against the abort rate $p_{\mathrm{abort}}$ at $p=0.001$, comparing Strategy 1 based on $Q_{\mathrm{LLR}}^{(2)}$ with the heuristic baselines and a gap proxy ($N_{\text{class}}=18$, “random” heuristic). Shaded regions represent 95% confidence intervals. (b) Various strategies are compared at a fixed $p_{\mathrm{abort}}=0.3$.
@@
We evaluate our approach on a $(3,4)$-regular HGP code with parameters $[[225,9,6]]$, constructed by taking the product of a 12-variable classical LDPC code with itself, whose check matrix is given in the Methods section. This particular code offers practical advantages with its depth-8 syndrome extraction circuit [42], which we employ in our simulations. Figure 5 shows the results at $p=0.001$, where our cluster-based approach achieves approximately two orders of magnitude improvement in $p_{\mathrm{log}}$ compared to the heuristic baselines, which is consistent with the performance gains observed for BB codes. Additionally, the gap proxy with $N_{\mathrm{class}}=18$ does not outperform the correction weight, again confirming its ineffectiveness.
Additional simulation results are provided in the Supplementary Material:
  • Comprehensive data for all parameter combinations (Supplementary Figures 1–3).
  • A detailed examination of the influence of the norm order $\alpha$ (Supplementary Figure 4), showing that $\alpha=1$ or $2$ is optimal for most cases.
  • An evaluation of the alternative “conv-max-conf” setting, which treats BP-converged samples as having maximum decoding confidence (Supplementary Figure 5). This setting avoids the need to modify BP+LSD to enforce LSD execution regardless of BP convergence, but is found to degrade post-selection performance in most cases.
  • A comprehensive analysis of gap proxy heuristics and true logical gap (Supplementary Note 1, including Supplementary Table 1 and Supplementary Figures 6–8). We compare various heuristics for selecting which logical classes to decode, including uniform random sampling (used in Figs. 4 and 5), weighted sampling based on the empirical logical error distribution, and adaptive variants. Surprisingly, none of these heuristics outperform naive uniform random sampling. Through exhaustive comparative decoding across all $2^{12}=4096$ logical classes for the $[[144,12,12]]$ BB code, we explain this counterintuitive result: post-selection performance depends on the discrimination power between successful and failed samples, rather than how accurately the metric approximates the true logical gap. Distribution-based heuristics improve gap approximation primarily for successful shots but not for failed shots, thereby reducing discriminative power.
@@ 2.4.2 Real-time strategy analysis
We now analyze the real-time post-selection strategy. To fairly assess the impact of “aborting mid-way”, we evaluate the average time cost per accepted shot $\bar{T}_{\mathrm{accepted}}$ instead of $p_{\mathrm{abort}}$, by summing the total number of syndrome extraction rounds elapsed across all samples (including both aborted and accepted runs) and dividing by the number of accepted samples. In other words, $\bar{T}_{\mathrm{accepted}}$ represents the expected time cost required to obtain a single accepted run when immediately retrying after each aborted run. For comparison, global strategies yield $\bar{T}_{\mathrm{accepted}}=T/(1-p_{\mathrm{abort}})$.
@@
Our analysis focuses on $T$-round memory experiments using the surface code with distance $d=13$ and the $[[144,12,12]]$ bivariate bicycle (BB) code. We assume the same circuit-level noise model as the global strategy analysis with $p\in\{0.003,0.005\}$. The decoding is performed using the sliding window framework with window parameters $(W,F)=(5,1)$ for the surface code and $(W,F)=(3,1)$ for the BB code. In both cases, the BP+LSD decoder serves as the inner decoder. For post-selection, we employ the cluster LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$ as our confidence metric with the lookback window size $L\in\{1,2,3,5,7\}$. By varying the cutoff value for this metric, we obtain different trade-offs between $p_{\mathrm{log}}$ and the average time cost per accepted shot $\bar{T}_{\mathrm{accepted}}$.
@@
Figure 6 presents our main results for both codes with $T=d$, clearly exhibiting the trade-off relation between $p_{\mathrm{log}}$ and $\bar{T}_{\mathrm{accepted}}$. For comparison, we include two baseline results: the global post-selection strategy (dashed lines) and standard sliding-window decoding without post-selection ($\times$” marks). Notably, our real-time strategy achieves performance comparable to the global strategy in most cases, and specifically for the BB code at $p=0.005$, it even surpasses global post-selection performance by at most around 1.5 orders of magnitude in $p_{\mathrm{log}}$.
@@
Figure 6: Analysis of real-time post-selection via Strategy 2 for (a) the surface code with $d=13$ and (b) the $[[144,12,12]]$ bivariate bicycle code. Logical error rates $p_{\mathrm{log}}$ are plotted against the average time cost per accepted shots $\bar{T}_{\mathrm{accepted}}$, which represents the average time cost required to succeed when immediately retrying after each aborted attempt. The results are for the memory experiments with $T=d$ at $p\in\{0.003,0.005\}$, decoded with the $(5,1)$ or $(3,1)$ sliding window method. The LLR 2-norm fraction $Q_{\mathrm{LLR}}^{(2)}$ is used for the metric and the parameter $L$ of the strategy varies across $\{1,2,3,5,7\}$. For comparison, the values for the global strategy (Strategy 1) and those for sliding window decoding without post-selection are presented additionally as dashed lines and $\times$” marks, respectively. Shaded regions represent 95% confidence intervals.
@@
The real-time strategy can be particularly useful for circuits involving many syndrome extraction rounds, where the global strategy may become computationally prohibitive. To assess the practical viability of our approach in such scenarios, we investigate how its performance scales with the total number of rounds $T$. Figure 7(a) examines the scaling behavior by plotting both $p_{\mathrm{log}}/T$ and the acceptance rate $1-p_{\mathrm{abort}}$ (in a logarithmic scale) against $T$ for the $[[144,12,12]]$ BB code under various fixed cutoff values, including the scenario without post-selection as a baseline. (Note that we here plot the acceptance rate rather than $\bar{T}_{\mathrm{accepted}}$, as it directly reveals its decay over rounds and allows us to estimate the per-round abort rate.)
Two key observations emerge from Fig. 7(a): First, $p_{\mathrm{log}}/T$ remains nearly constant across different $T$ values even under post-selection, which indicates that logical errors accumulate linearly with the number of rounds, preserving the expected scaling behavior shown in the case without post-selection. Second, $\log(1-p_{\mathrm{abort}})$ exhibits clear linear dependence on $T$. This behavior can be interpreted as every single round has an equal per-round abort rate $p_{\mathrm{abort}}^{\mathrm{round}}$, which is related to $p_{\mathrm{abort}}$ as $p_{\mathrm{abort}}=1-(1-p_{\mathrm{abort}}^{\mathrm{round}})^{T}$ and thus can be estimated from the slope of the fit in the right panel. In Fig. 7(b), we plot the estimated values of $p_{\mathrm{log}}/T$ (at $T=24$) and $p_{\mathrm{abort}}^{\mathrm{round}}$ against the cutoff, clearly illustrating the trade-off relation between them. Together, these properties ensure that the real-time strategy offers predictable scaling behavior, allowing reliable extrapolation of performance to arbitrary number of rounds.